An electromagnet exerts a force of 15 N and moves a soft iron armature through a distance of 12 mm in 50ms. Determine the power consumed


1
Expert's answer
2022-06-07T17:37:29-0400

Given:

"F=15N;"

"d=12mm=0.012m;"

"t=50ms=0.05s."

Find:

"P-?"

Solution:

Power consumed = (Work done by electromagnetic force)/(time);

"P=\frac{W}{t};" "W=Fdcos\alpha" ; here "\alpha=0;" So,

"P=\frac{Fd}{t}=\frac{15N\times0.012m}{0.05s}=3.6W."

Answer:

"P=3.6W."